# Connected Components in Linear Work and Near-Optimal Time

### v2

@@ 1 Introduction
For a long time, an interesting question is whether we can beat the $O(\log n)$-time barrier for computing connected components in parallel. In this respect, the initial breakthrough was made in the Massively Parallel Computation (MPC) model [BKS17], a model that is stronger than the classic PRAM model [KSV10, GSZ11, ASW19]. The work of [${}^{+}$18] showed an algorithm that completes in $O(\log d\cdot\log\log_{m/n}n)$ rounds of the MPC model, where $d$ denotes the maximum diameter over all components of the graph. Their result shows that for graphs with small diameter, we can indeed overcome the $\log n$ barrier, at least in more powerful parallel computation models such as the MPC model. The subsequent work of [${}^{+}$19] further improved the result, by presenting an MPC algorithm that completes in $O(\log d+\log\log_{m/n}n)$ rounds. This line of work culminated in the recent work of Liu, Tarjan, and Zhong [LTZ20]. Inspired by the MPC algorithms, [LTZ20] showed that even on a PRAM, one can compute connectivity in $O(\log d+\log\log_{m/n}n)$ parallel time. Note that accomplishing the result on a PRAM is much more non-trivial than that in the MPC model, since in the MPC model, primitives such as sorting and prefix sum can be accomplished in constant rounds, whereas on a CRCW PRAM, these primitives require $\Omega(\log n/\log\log n)$ time even using $\text{poly}(n)$ processors [BH89].
@@ 1.1 Our Results and Contributions
First, we devise a linear-work algorithm that contracts the graph to $n/\beta$ vertices in $O(\log\beta)$ time for any $\beta\geq\log n$. It is well-known that random leader election contracts the graph to $n/\beta$ vertices in $O(\log\beta)$ time for any $\beta>0$, but it requires either an MPC (see [${}^{+}$19] and the references therein) or PRAM with super-linear work [Rei84, Gaz91]. Note that for existing linear-work algorithms, contracting to $n/\beta$ vertices requires $\Omega((\log\beta)^{2})$ time and only works for $\beta\geq\log n$ (see §3 for details). Graph contraction is a fundamental graph primitive and we hope our new algorithm can find more applications.
Second, we devise a new technique that densifies the graph without decreasing the component-wise spectral gap, and the total work is linear in the number of edges in the original graph. Graph densification is used in [${}^{+}$18, ${}^{+}$19, ${}^{+}$19, LTZ20, ${}^{+}$23] to achieve $o(\log n)$ running time, but they either take super-linear work (on a PRAM) or can decrease the spectral gap of the graph (by adding extra edges to the graph). We hope our technique of graph densification that preserves the component-wise spectral gap can find further applications in graph algorithms and spectral graph theory.
On the lower bound side, the popular 2-Cycle Conjecture [RVW18, ASW19, ${}^{+}$18, ${}^{+}$19] states that any MPC algorithm with $n^{1-\Omega(1)}$ memory per processor cannot distinguish between one cycle of size $n$ with two cycles of size $n/2$ in $o(\log n)$ rounds. Based on this, [${}^{+}$19] shows that any MPC algorithm in this setting requires $\Omega(\log d)$ rounds to compute connected components of any given graph with diameter $d\geq\operatorname{poly}(\log(n))$ with high probability. In this paper, we use a similar approach to show that any PRAM algorithm with total memory $O(m+n)$ requires $\Omega(\log(1/\lambda))$ time to compute connected components of any given graph with minimum spectral gap $\lambda\leq 1/\operatorname{poly}(\log(n))$ with high probability, unless the 2-Cycle Conjecture is wrong. This shows that our algorithm is conditional optimal in the running time as a parameter of $\lambda$.
@@ 1.2 Additional Related Work
It is also interesting to compare our result with the earlier work of Assadi, Sun, and Weinsten [ASW19], who showed an MPC algorithm for computing connectivity in $O(\log(1/\lambda)+\log\log n)$ rounds when $\lambda$ is known to the algorithm beforehand. When their algorithm has no prior knowledge on $\lambda$, the running time degenerates to $O(\log(1/\lambda)+\log\log(1/\lambda)\cdot\log\log n)$ and requires slightly more processors. Our connectivity algorithm does not require prior knowledge on $\lambda$ and achieves the same round complexity on a PRAM. As mentioned, the PRAM model is less powerful than MPC and thus our result requires more non-trivial techniques than that of [ASW19]. Indeed, our techniques have no overlaps with that of [ASW19]. In [ASW19], they have two novel ideas. First they show that in the MPC model, independent random walks for all vertices can be found using only small number of rounds. The other idea used in the paper is an $O(\log\log n)$ round algorithm for finding spanning forests in random graphs. They provide a method to transform each connected component of the input graph into a random graph, and later use the algorithm for random graphs to find connected components.
Coy and Czumaj [CC21] showed how to make the MPC algorithms of [${}^{+}$18, ${}^{+}$19] deterministic, while preserving the $O(\log d+\log\log_{m/n}n)$ round complexity. One interesting future direction is how to deterministically compute connectivity in sub-logarithmic parallel time on a PRAM with $O(m+n)$ processors. Note that for this question, no answer is known even if we are willing to tolerate sub-optimality in total work.
In a very recent work, [${}^{+}$23] showed a deterministic MPC algorithm for computing connected components of any forest in $O(\log d)$ rounds.
@@ 2.1 Definitions and Notations
We assume the ARBITRARY CRCW PRAM model [Vis83], where each processor has constant number of cells (i.e., words) in its private memory. All processors share a common memory array. The processors run synchronously. In one step, each processor can read a cell in memory, write to a cell in memory, and do a constant amount of local computation. Any number of processors can read from or write to the same common memory cell concurrently. If more than one processors write to the same memory cell at the same time, an arbitrary one succeeds, and the algorithm should be correct no matter which one succeeds.
@@
Given graph $G=(V,E)$, a contraction on $G$ is formed by identifying two distinct vertices $u,v\in V$ into a single vertex $v^{*}$, i.e., replace vertices $u,v$ with $v^{*}$ in $V$, and replace any edge $(u,w),(v,w)$ with $(v^{*},w)$ in $E$ for all $w$. A contraction algorithm is an algorithm that performs a set of contractions on the input graph. To get a correct connected component algorithm, each contraction is limited to vertices from the same component. If $f,g$ are both contraction algorithms, then $f\circ g$ is also a contraction algorithm by the union of their sets of contractions.
@@ 2.2 Preliminaries on Spectral Graph Theory
[…]
@@ 3 Technical Overview
One of our key ideas is to devise a new method to sparsify the graph to only $O(m/\operatorname{poly}(\log n))$ edges without affecting connectivity. To achieve this, our algorithm first contracts and consequently densifies the graph to make each vertex (except those in very small connected components) sufficiently large (vertex-) degree. At this moment, we then use random edge sampling to sparsify the graph without breaking the connectivity of each connected component w.h.p. (provably). For comparison and highlight, recall that in [${}^{+}$18] (and also in [LTZ20]), they also densify the graph to minimum degree at least $b$ such that each vertex contracts to a leader w.h.p. after sampling each vertex as a leader w.p. $(\log n)/b$. However, their densification is by adding extra edges to the graph, which might decrease the spectral gap of a component and thus increase the running time of our algorithm. (Our densification is by contractions so it preserves the spectral gaps of all components.) More importantly, their densification requires $\Theta(m\log d)$ work since they need to use all the $m$ edges and repeat the broadcasting for $\Theta(\log d)$ times; by contrast, ours only takes $O(m)$ work.
@@
Connectivity with known $\lambda\geq 1/\log n$ (the assumption on $\lambda$ will be removed in §3.4): Stage $1$: Contract the graph to $n/(\log n)^{20}$ vertices and assign each vertex $(\log n)^{20}$ processors. Stage $2$: Increase the degree of each vertex to $(\log n)^{5}$ using the $(\log n)^{20}$ processors per vertex. Stage $3$: Sample each edge w.p. $1/(\log n)$ then compute the connected components of the sampled subgraph.
@@ 3.1 Stage $1$: Contract the Graph to Reduce Vertices
In Stage $1$ we first present a linear-work algorithm that reduces the number of vertices in the graph to $n/\log\log n$ in $O(\operatorname{poly}(\log\log\log n))$ time. To achieve this reduction in the number of vertices, we borrow some ideas from Gazit’s paper. The main tool that we use is a constant-time and linear-work algorithm that reduces the number of vertices by a constant factor. Our constant-shrink algorithm is inspired by the matching MPC algorithm of [${}^{+}$19]. In each call, our algorithm finds a large matching between the vertices of the graph such that at least a constant fraction of the vertices are matched. We can then contract these matching edges to reduce the number of vertices by a constant factor in each call.
@@
On the contrary, we only repeat the above separation process for $O(\log\log\log n)$ times and each of them takes $O(\log\log\log n)$ time. One can show that the number of vertices with high degree in the original graph is reduced to $n/(\log n)^{21}$ with high probability. We prove that the sparse part has $O(n/\log\log n)$ edges with high probability. This is done by showing that each high-degree vertex contributes no edge to the sparse part with high probability, and the total contribution of edges from low-degree vertices is $O(n/\log\log n)$ with high probability by the fact that the number of vertices is at most $n/\log\log n$ and with the use of concentration bounds for self-bounding functions. So we can contract the sparse part to at most $n/(\log n)^{21}$ vertices by $O(\log\log n)$ applications of the constant-shrink algorithm in $O(n)$ work since the number of edges is bounded from above by $O(n/\log n)$ with high probability.
@@ 3.2 Stage $2$: Increase the Minimum Degree
We observe that the potential argument in [${}^{+}$19] (and thus [LTZ20]) implies that after $\Theta(\log r+\log\log n)$ rounds of the algorithm, each vertex has edges to all vertices within distance $2^{r}$ in the original graph. This is because each vertex either increases its level or connects to all vertices within distance $2$ (for simplicity, we ignore the issue of altering graph edges). This implies an algorithm that increases the minimum degree of the graph to $(\log n)^{5}$ in $O(\log\log n)$ time. However, such an algorithm requires $\Omega(m\log\log n)$ work since it uses all $m$ edges in each round. Instead, we run $O(\log\log n)$ rounds of the algorithm in [LTZ20] on the skeleton graph described above. If the component has small size, then it is the same component as in the original graph and we contract it to $1$ vertex. Otherwise the component has size at least $(\log n)^{10}$ and there are two cases. If the component has diameter at most $(\log n)^{10}$ then calling [LTZ20] computes this component. This component can be different from the component in the input graph but the contracted vertex has at least $(\log n)^{10}$ children, giving at least $(\log n)^{10}/2$ adjacent edges in the contracted graph. Otherwise, there exists a shortest path of length at least $(\log n)^{10}$, and each vertex ends up with at least $(\log n)^{5}$ new neighbors by the potential argument, giving a contracted graph with minimum degree at least $(\log n)^{5}$.
@@ 3.4 Removing the Assumption on Spectral Gap
It is easy to see that $O(\log\log n)$ phases are sufficient to obtain a correct connected component algorithm, because any graph with $m\leq n^{c}$ has minimum spectral gap $\lambda\geq(1/\log n)^{(3/2)^{9c\log\log n}}$.
@@
Connectivity with unknown $\lambda$: Parameter $b=(\log n)^{20}$. Build random subgraph $H^{\prime},H^{\prime\prime}$ by sampling each edge in the original graph $G$ w.p. $1/(\log n)^{3}$. For phase $i$ from 00 to $O(\log\log n)$: Run $O(\log b)$ steps of Stage $1$ to decrease the number of vertices to $O(n/b)$. To build $H$, preserve all edges in $G$ adjacent to vertices with degree at most $b$ in $H^{\prime}$, then add all edges of $H^{\prime}$ into $H$. Increase the minimum degree of $H$ to $b$ by contraction. Perform contractions on $H^{\prime\prime}$ as performed on $H$ in the previous step. Try to compute the connected components of $H^{\prime\prime}$ in $O(\log b)$ time. If all connected components of $H^{\prime\prime}$ are computed then return. Else update $b\leftarrow b^{3/2}$.
@@ 4 Stage 1: Contracting the Graph to $n/\operatorname{poly}(\log n)$ Vertices
We begin with presenting a linear-work algorithm that reduces the number of vertices in the graph to $n/\log\log n$ in $O((\log\log\log n)^{2})$ time. The main tool that we use is a constant-time and linear-work algorithm that reduces the number of vertices by a constant factor. Our constant-shrink algorithm is inspired by the matching MPC algorithm of [${}^{+}$19]. In each call, our algorithm finds a large matching between the vertices of the graph such that at least a constant fraction of the vertices are matched. We can then contract these matching edges to reduce the number of vertices by a constant factor in each call.

### v4

@@ 1 Introduction
For a long time, an interesting question is whether we can beat the $O(\log n)$-time barrier for computing connected components in parallel. In this respect, the initial breakthrough was made in the Massively Parallel Computation (MPC) model [BKS17], a model that is stronger than the classic PRAM model [KSV10, GSZ11, ASW19]. The work of [ASS+18] showed an algorithm that completes in $O(\log d\cdot\log\log_{m/n}n)$ rounds of the MPC model, where $d$ denotes the maximum diameter over all components of the graph. Their result shows that for graphs with small diameter, we can indeed overcome the $\log n$ barrier, at least in more powerful parallel computation models such as the MPC model. The subsequent work of [BDE+19] further improved the result, by presenting an MPC algorithm that completes in $O(\log d+\log\log_{m/n}n)$ rounds. This line of work culminated in the recent work of Liu, Tarjan, and Zhong [LTZ20]. Inspired by the MPC algorithms, [LTZ20] showed that even on a PRAM, one can compute connectivity in $O(\log d+\log\log_{m/n}n)$ parallel time. Note that accomplishing the result on a PRAM is much more non-trivial than that in the MPC model, since in the MPC model, primitives such as sorting and prefix sum can be accomplished in constant rounds, whereas on a CRCW PRAM, these primitives require $\Omega(\log n/\log\log n)$ time even using $\text{poly}(n)$ processors [BH89].
@@ 1.1 Our Results and Contributions
First, we devise a linear-work algorithm that contracts the graph to $n/\beta$ vertices in $O(\log\beta)$ time for any $\beta\geq\log n$. It is well-known that random leader election contracts the graph to $n/\beta$ vertices in $O(\log\beta)$ time for any $\beta>0$, but it requires either an MPC (see [BDE+19] and the references therein) or PRAM with super-linear work [Rei84, Gaz91]. Note that for existing linear-work algorithms, contracting to $n/\beta$ vertices requires $\Omega((\log\beta)^{2})$ time and only works for $\beta\geq\log n$ (see §3 for details). Graph contraction is a fundamental graph primitive and we hope our new algorithm can find more applications.
Second, we devise a new technique that densifies the graph without decreasing the component-wise spectral gap, and the total work is linear in the number of edges in the original graph. Graph densification is used in [ASS+18, BDE+19, LTZ20, BLM+23] to achieve $o(\log n)$ running time, but they either take super-linear work (on a PRAM) or can decrease the spectral gap of the graph (by adding extra edges to the graph). We hope our technique of graph densification that preserves the component-wise spectral gap can find further applications in graph algorithms and spectral graph theory.
On the lower bound side, the popular 2-Cycle Conjecture [RVW18, ASW19, ASS+18, BDE+19] states that any MPC algorithm with $n^{1-\Omega(1)}$ memory per processor cannot distinguish between one cycle of size $n$ with two cycles of size $n/2$ in $o(\log n)$ rounds. Based on this, [BDE+19] shows that any MPC algorithm in this setting requires $\Omega(\log d)$ rounds to compute connected components of any given graph with diameter $d\geq\operatorname{poly}(\log(n))$ with high probability. In this paper, we use a similar approach to show that any PRAM algorithm with total memory $O(m+n)$ requires $\Omega(\log(1/\lambda))$ time to compute connected components of any given graph with minimum spectral gap $\lambda\leq 1/\operatorname{poly}(\log(n))$ with high probability, unless the 2-Cycle Conjecture is wrong. This shows that our algorithm is conditional optimal in the running time as a parameter of $\lambda$.
@@ 1.2 Additional Related Work
It is also interesting to compare our result with the earlier work of Assadi, Sun, and Weinsten [ASW19], who showed an MPC algorithm for computing connectivity in $O(\log(1/\lambda)+\log\log n)$ rounds when $\lambda$ is known to the algorithm beforehand. When their algorithm has no prior knowledge on $\lambda$, the running time degenerates to $O(\log(1/\lambda)+\log\log(1/\lambda)\cdot\log\log n)$ and requires slightly more processors. Our connectivity algorithm does not require prior knowledge on $\lambda$ and achieves the same round complexity on a PRAM. As mentioned, the PRAM model is less powerful than MPC and thus our result requires more non-trivial techniques than those of [ASW19]. Indeed, our techniques have no overlaps with those of [ASW19]. In [ASW19], they have two novel ideas. First they show that in the MPC model, independent random walks for all vertices can be found using only small number of rounds. The other idea used in the paper is an $O(\log\log n)$ round algorithm for finding spanning forests in random graphs. They provide a method to transform each connected component of the input graph into a random graph, and later use the algorithm for random graphs to find connected components.
Coy and Czumaj [CC22] showed how to make the MPC algorithms of [ASS+18, BDE+19] deterministic, while preserving the $O(\log d+\log\log_{m/n}n)$ round complexity. One interesting future direction is how to deterministically compute connectivity in sub-logarithmic parallel time on a PRAM with $O(m+n)$ processors. Note that for this question, no answer is known even if we are willing to tolerate sub-optimality in total work.
In a very recent work, [BLM+23] showed a deterministic MPC algorithm for computing connected components of any forest in $O(\log d)$ rounds.
@@ 2.1 Definitions and Notations
We assume the ARBITRARY CRCW PRAM model [Vis83], where each processor has constant number of cells (i.e., words) in its private memory. All processors share a common memory array. The processors run synchronously. In one step, each processor can read a cell in memory, write to a cell in memory, and do a constant amount of local computation. Any number of processors can read from or write to the same common memory cell concurrently. If more than one processor write to the same memory cell at the same time, an arbitrary one succeeds, and the algorithm should be correct no matter which one succeeds.
@@
Given graph $G=(V,E)$, a contraction on $G$ is formed by identifying two distinct vertices $u,v\in V$ into a single vertex $v^{*}$, i.e., replace vertices $u,v$ with $v^{*}$ in $V$, and replace any edge $(u,w),(v,w)$ with $(v^{*},w)$ in $E$ for all $w$. A contraction algorithm is an algorithm that performs a set of contractions on the input graph. To get a correct connected components algorithm, each contraction is limited to vertices from the same component. If $f,g$ are both contraction algorithms, then $f\circ g$ is also a contraction algorithm by the union of their sets of contractions.
@@ 2.2 Preliminaries on Spectral Graph Theory
$\mathcal{L}(u,v)=\begin{cases}1-\frac{w(v,v)}{\deg(v)}&\text{if $u=v$, and $\deg(v)\neq 0$}\\ -\frac{w(u,v)}{\sqrt{\deg(u)\deg(v)}}&\text{if $u$ and $v$ are adjacent}\\ 0&\text{otherwise}.\end{cases}$
@@ 3 Technical Overview
One of our key ideas is to devise a new method to sparsify the graph to only $O(m/\operatorname{poly}(\log n))$ edges without affecting connectivity. To achieve this, our algorithm first contracts and consequently densifies the graph to make each vertex (except those in very small connected components) sufficiently large (vertex-) degree. At this moment, we then use random edge sampling to sparsify the graph without breaking the connectivity of each connected component w.h.p. (provably). For comparison and highlight, recall that in [ASS+18] (and also in [LTZ20]), they also densify the graph to minimum degree at least $b$ such that each vertex contracts to a leader w.h.p. after sampling each vertex as a leader w.p. $(\log n)/b$. However, their densification is by adding extra edges to the graph, which might decrease the spectral gap of a component and thus increase the running time of our algorithm. (Our densification is by contractions so it preserves the spectral gaps of all components.) More importantly, their densification requires $\Theta(m\log d)$ work since they need to use all the $m$ edges and repeat the broadcasting for $\Theta(\log d)$ times; by contrast, ours only takes $O(m)$ work.
@@
Connectivity with known $\lambda\geq 1/\log n$ (the assumption on $\lambda$ will be removed in §3.4):
  • Stage $1$: Contract the graph to $n/(\log n)^{20}$ vertices and assign each vertex $(\log n)^{20}$ processors.
  • Stage $2$: Increase the degree of each vertex to $(\log n)^{5}$ using the $(\log n)^{20}$ processors per vertex.
  • Stage $3$: Sample each edge w.p. $1/(\log n)$ then compute the connected components of the sampled subgraph.
@@ 3.1 Stage $1$: Contract the Graph to Reduce Vertices
In Stage $1$ we first present a linear-work algorithm that reduces the number of vertices in the graph to $n/\log\log n$ in $O(\operatorname{poly}(\log\log\log n))$ time. To achieve this reduction in the number of vertices, we borrow some ideas from Gazit’s paper. The main tool that we use is a constant-time and linear-work algorithm that reduces the number of vertices by a constant factor. Our constant-shrink algorithm is inspired by the matching MPC algorithm of [BDE+19]. In each call, our algorithm finds a large matching between the vertices of the graph such that at least a constant fraction of the vertices are matched. We can then contract these matching edges to reduce the number of vertices by a constant factor in each call.
@@
On the contrary, we only repeat the above separation process for $O(\log\log\log n)$ times and each of them takes $O(\log\log\log n)$ time. One can show that the number of vertices with high degree in the original graph is reduced to $n/(\log n)^{21}$ with high probability. We prove that the sparse part has $O(n/\log\log n)$ edges with high probability. This is done by showing that each high-degree vertex contributes no edge to the sparse part with high probability, and the total contribution of edges from low-degree vertices is $O(n/\log\log n)$ with high probability by the fact that the number of vertices is at most $n/\log\log n$ and with the use of concentration bounds for self-bounding functions. So we can contract the sparse part to at most $n/(\log n)^{21}$ vertices by $O(\log\log n)$ applications of the constant-shrink algorithm in $O(n)$ work since the number of edges is bounded from above by $O(n/\log\log n)$ with high probability.
@@ 3.2 Stage $2$: Increase the Minimum Degree
We observe that the potential argument in [BDE+19] (and thus [LTZ20]) implies that after $\Theta(\log r+\log\log n)$ rounds of the algorithm, each vertex has edges to all vertices within distance $2^{r}$ in the original graph. This is because each vertex either increases its level or connects to all vertices within distance $2$ (for simplicity, we ignore the issue of altering graph edges). This implies an algorithm that increases the minimum degree of the graph to $(\log n)^{5}$ in $O(\log\log n)$ time. However, such an algorithm requires $\Omega(m\log\log n)$ work since it uses all $m$ edges in each round. Instead, we run $O(\log\log n)$ rounds of the algorithm in [LTZ20] on the skeleton graph described above. If the component has small size, then it is the same component as in the original graph and we contract it to $1$ vertex. Otherwise the component has size at least $(\log n)^{10}$ and there are two cases. If the component has diameter at most $(\log n)^{10}$ then calling [LTZ20] computes this component. This component can be different from the component in the input graph but the contracted vertex has at least $(\log n)^{10}$ children, giving at least $(\log n)^{10}/2$ adjacent edges in the contracted graph. Otherwise, there exists a shortest path of length at least $(\log n)^{10}$, and each vertex ends up with at least $(\log n)^{5}$ new neighbors by the potential argument, giving a contracted graph with minimum degree at least $(\log n)^{5}$.
@@ 3.4 Removing the Assumption on Spectral Gap
It is easy to see that $O(\log\log n)$ phases are sufficient to obtain a correct connected components algorithm, because any graph with $m\leq n^{c}$ has minimum spectral gap $\lambda\geq(1/\log n)^{(3/2)^{9c\log\log n}}$.
@@
Connectivity with unknown $\lambda$:
  • Parameter $b=(\log n)^{20}$.
  • Build random subgraph $H^{\prime},H^{\prime\prime}$ by sampling each edge in the original graph $G$ w.p. $1/(\log n)^{3}$.
  • For phase $i$ from 00 to $O(\log\log n)$:
    • Run $O(\log b)$ steps of Stage $1$ to decrease the number of vertices to $O(n/b)$.
    • To build $H$, preserve all edges in $G$ adjacent to vertices with degree at most $b$ in $H^{\prime}$, then add all edges of $H^{\prime}$ into $H$.
    • Increase the minimum degree of $H$ to $b$ by contraction.
    • Perform contractions on $H^{\prime\prime}$ as performed on $H$ in the previous step.
    • Try to compute the connected components of $H^{\prime\prime}$ in $O(\log b)$ time.
    • If all connected components of $H^{\prime\prime}$ are computed then return.
    • Else update $b\leftarrow b^{3/2}$.
@@ 4 Stage 1: Contracting the Graph to $n/\operatorname{poly}(\log n)$ Vertices
We begin with presenting a linear-work algorithm that reduces the number of vertices in the graph to $n/\log\log n$ in $O((\log\log\log n)^{2})$ time. The main tool that we use is a constant-time and linear-work algorithm that reduces the number of vertices by a constant factor. Our constant-shrink algorithm is inspired by the matching MPC algorithm of [BDE+19]. In each call, our algorithm finds a large matching between the vertices of the graph such that at least a constant fraction of the vertices are matched. We can then contract these matching edges to reduce the number of vertices by a constant factor in each call.